OpenStudy (brooke1556):

Write the equation of the line that passes through (3, –2) and has a slope of 4 in point-slope form. y + 2 = 4(x – 3) y – 3 = 4(x + 2) x – 3 = 4(y + 2) x + 2 = 4(y – 3)

OpenStudy (mrnood):

OK - this is just a case of putting numbers into the formula It tells you the slope is 4 and the point (x1,y1) is (3,-2) the equation is y -y1 = m(x-x1) just put the numbers in (hint: be careful of the minus signs....)

OpenStudy (mrnood):

btw : m is the slope (=4 in this case)

OpenStudy (brooke1556):

So 3 - (-2) = 5?

OpenStudy (mrnood):

look at the format of teh answers - they have x and y in them the answer looks EXACTLY like this: y -y1 = m(x-x1) but you need to put in the values for y1, m and x1 so write the equation in that form using the point and slope you were given in the question

OpenStudy (brooke1556):

Oh okay, So would it be, y+2=4(x-3) ?

OpenStudy (mrnood):

Yes - well done

OpenStudy (brooke1556):

Thank you!